# Strain effects in a directly bonded diamond-on-insulator substrate

### v1

@@ 1 Introduction
One of the challenges in achieving robust quantum communication and high-fidelity quantum computing has long been the entanglement generation rate between NV centers, which is limited by the relatively low fraction of emitted coherent photons that can be detected. A solution to address this problem is to integrate NV centers into optical cavities with a high ratio of quality factor versus mode volume. This is done to increase the zero-phonon line (ZPL) emission and improve collection efficiency [1].
@@
However, despite the effectiveness of direct bonding and subsequent annealing in enhancing interface quality, this fabrication process can introduce unintended strain effects within the diamond structure. It is very likely that the diamond, having undergone the aforementioned chemical processing, will exhibit further stress effects on the surface which is bonded to the insulator substrate. Furthermore, part of the direct bonding process takes place under a temperature of 200°C. During the cooling period, each of the bonded materials contracts at different rates, according to their unique thermal expansion coefficients. This mismatch between the thermal expansion rates between the diamond and the silicon dioxide/silicon can lead to the generation of strain at their interface [14, 15], and potentially further deep into the diamond lattice.
These strain effects can cause deformations to the crystal lattice, which result in both inhomogeneous broadening and degrade the spin dephasing time of the NV centers. Therefore, it is deemed critical to be able to understand and visually map the effects of strain, providing meaningful insight into the quality of the bonded samples.
@@ 2 Direct bonding process overview
Hydrophilic direct bonding is a process that involves creating a bond between surfaces treated to exhibit hydrophilic properties. The first step is surface activation, which prepares the surfaces for subsequent bonding by introducing reactive hydroxyl groups (–OH). For the diamond substrate, this involves thorough cleaning using a Piranha solution, a highly oxidative mixture of sulfuric acid (H2SO4) and hydrogen peroxide (H2O2) in a 3:1 ratio. This treatment results in the termination of the diamond surface with hydroxyl groups (C–OH), enhancing its reactivity, by providing essential reactive sites for bonding with other molecules, or facilitating further chemical reactions [13, 16].
The insulator substrate is composed of a layer of SiO2 on Si. This SiO2/Si substrate undergoes O2 plasma treatment, a common method for cleaning and oxidizing silicon surfaces. This process forms silanol groups Si-OH on the substrate’s surface, which promote adhesion, modify the surface energy, and enable further chemical reactions or the attachment of functional molecules [13, 16].
Once the surfaces are activated, the initial bond formation between the diamond and SiO2/Si substrate occurs through van der Waals forces and hydrogen bonding, facilitated by the presence of water molecules. Post-activation, both surfaces become highly hydrophilic, due to the presence of hydroxyl groups (-OH). This hydrophilicity causes the surfaces to attract and adsorb water molecules from the environment, forming a thin, hydrated layer on each surface. Van der Waals forces and hydrogen bonding bring the surfaces close together and stabilize their initial contact. The hydroxyl groups (C–OH and Si-OH) and the adsorbed water molecules create hydrogen bonds between them, that facilitate alignment and bonding of the surfaces. Hydrogen bonds, although individually weak, contribute significantly to the initial adhesion between the surfaces [17].
@@
After the initial bond formation, the bonded DOI substrate undergoes an annealing process. The annealing temperature and duration are critical parameters for achieving optimal bond strength. During annealing, the bond strength is enhanced through a dehydration reaction [17], where water molecules gradually diffuse through the interface and form a strong covalent bond (C–O–Si), expressed by the reaction seen below:
C-OH + HO-Si → C-O-Si + H2O
Table 1 below summarizes the hydrophilic direct bonding process steps that are performed in this research.
Table 1: Process steps for hydrophilic direct bonding of a DOI substrate
Process Step Parameters / Remarks
Piranha treatment Immerse diamond in a solution of 30 ml H2SO4 (95–97%) and 10 ml H2O2 (31%) for 30 minutes at 75°C.
Water rinsing of diamond Rinse with deionized (DI) water for 5–10 minutes. Do not dry with N2 gas.
SiO2 PECVD Deposit SiO2 on Si wafer using PECVD, then dice into $25\times 25$ mm2 chips.
Water rinsing of SiO2/Si substrate Rinse with DI water for 5 minutes and dry with N2 gas.
O2 plasma activation Oxidize using PVA Tepla tool at 1000 W and O2 flow of 400 ml/min for 5 minutes.
Contacting Bring diamond and SiO2/Si into contact at 20°C and 40% humidity.
Annealing Anneal at 200°C for 24 hours with N2 flow.
3 Strain in NV centers
3.1 Causes of strain in directly bonded DOI substrates
@@
The annealing process, however, is critical for enhancing the bond strength of the two different substrates. This is achieved through dehydration reactions, which further exacerbate the strain effects. During the dehydration of the interface, water molecules diffuse out and covalent bonds are formed between the diamond and the insulator substrate, leading to a slight atomic rearrangement. As a result, this lattice misalignment contributes to the strain effects in the diamond layer[20]. Moreover, the surface activation steps may contribute to strain effects. During the Piranha or the $\mathrm{O_{2}}$ plasma treatment, unwanted surface defects can be introduced, which may alter the surface properties when bonding takes place, leading to additional strain[21].
@@ 3.2.1 Spin-strain interaction
In reality, numerous imperfections exist within the diamond lattice, making the spin-strain interaction non-negligible. According to Udvarhelyi et al (2018) [23], the spin-strain interaction for the ground-state spin triplet can be described as a deformation in the diamond crystal, causing the strain tensor $\varepsilon_{ij}$ to couple to the NV spin through the symmetry-allowed Hamiltonian:
@@
In the NV reference frame $(x,y,z)$, aligned such that $z$ is the NV-axis and $(x,y)$ spans the perpendicular plane, the general spin-strain Hamiltonian allowed by the $C_{3v}$ symmetry is written as [23]:
@@
Table 2: Spin-strain ($h$) coupling-strength parameters calculated from density functional theory by Udvarhelyi et al (2018) [23].
@@ 3.2.2 Eigenfrequencies including the Zeeman term
$\frac{\hat{H}_{B}}{h}=\gamma\mathit{B}\cdot S,$ (5)
where $\gamma$ is the electron gyromagnetic ratio, $2.8$ MHz/G. After diagonalizing this matrix and including the spin-strain interaction Hamiltonian, the following equation is obtained:
@@
where $M_{x}$ and $M_{y}$ represent the shear (transverse) strain components, corresponding to lattice distortions perpendicular to the NV axis. These transverse components primarily influence the magnitude of peak splitting in the ODMR spectra.
Therefore, the effects of strain can be directly characterized by ODMR [24], as lattice distortions lead to measurable frequency shifts and splittings of the spin resonance transitions.
@@ 4 Experimental Section
The primary objective of this study is to measure the effects of strain in diamonds that have undergone direct bonding with an insulator ($\mathrm{SiO_{2}/Si}$) substrate. To achieve this, NV centers are used due to their excellent properties that make them highly sensitive to external magnetic fields and strain fields. As already described above, strain can cause splitting and shifts in the energy levels of NV centers.
To enable these measurements, a confocal Optically Detected Magnetic Resonance (ODMR) setup is used. With ODMR, these changes in the energy states and spins can be quantified in the form of the presence of a magnetic, electric, or strain field, or temperature variations. The frequency shift $\Delta f_{\mathrm{shift}}$ and frequency splitting $\Delta f_{\mathrm{split}}$ are related to the strain tensor components $\epsilon_{\mathrm{ij}}$, through Equation 3.
@@
For the experiment, a $10\times 10\,\mathrm{mm^{2}}$ bulk diamond substrate (supplied by EDP), with a thickness of $\sim 300$ µm is investigated. Let this sample be referred to as $\mathrm{DS_{PB}}$. It is a CVD-grown substrate, cut along the {100} crystallographic plane, that contains randomly distributed ensembles of NV centers with a concentration of roughly 15 ppb. This sample underwent the direct bonding process, but due to its large surface area and surface roughness, the result was partial bonding of the DOI substrate, meaning that two regions of interest are present—both bonded and unbonded—on the same sample. Different strain fields may exist in these regions, since only part of its surface formed a strong bond with the insulator substrate underneath it. This bonded region has potentially experienced increased strain effects compared to the unbonded one. By comparing the ODMR peak splitting and frequency shift values at the top surface and at the bottom surface (diamond-$\mathrm{SiO_{2}/Si}$ interface), an overall increase in strain is observed in the interface.
These samples had a few microns etched away from their top surface with $\mathrm{O_{2}}$ plasma, in order to allow focusing on their back surface (diamond-$\mathrm{SiO_{2}/Si}$ interface). This challenge arose due to the objective lens’ small working distance of 400 µm and the 100 µm diameter of the microwave delivery wire, which is placed between the objective and the piezo stage. Although the refractive index of diamond, $n=2.4$, provides some leniency by modifying the effective distance of the objective-to-diamond, further etching of the surface ensures that proper focus on the back surface can be achieved and eliminates any worries of collision of the diamond with the objective during 2D scanning. It should be noted that the following measurements are performed in the presence of an external bias magnetic field $B_{z}=15.7$ G (or 43.9 MHz splitting) along the $z$-axis. A 532 nm green laser is used for optical excitation, at a power of 0.25 mW, and the emitted fluorescence is collected via a single-photon counting module (COUNT-10C-FC), coupled to a fiber.
@@ 5 Results and Discussion
Figure 3: ODMR 2D scans performed in a $60\times 60$ µm2 area, depicting the effects of the volumetric strain amplitude $M_{z}$. a) Top surface, b) Back surface.
Afterward, 2D ODMR maps were obtained, starting with a $60\times 60$ µm2 area at the top surface, and then focusing on the diamond plane close to the diamond-$\mathrm{SiO_{2}/Si}$ interface. Figure 3 represents volumetric strain variation, obtained by calculating the frequency shift $\Delta f_{\mathrm{shift}}$ of every point in the map, relative to the zero field splitting (ZFS) value of $D=$ 2.87 GHz, and solving for $M_{z}$, as seen in Equation 9.
The milder volumetric strain effects at the top surface are considered to be part of the intrinsic strain of the diamond. However, moving closer towards the back surface shows increased strain effects, which are assumed to have been caused by the thermal expansion mismatch that occurs during the direct bonding process.
Figure 4: ODMR 2D scans performed in a $60\times 60$ µm2 area, depicting the effects of the transverse strain amplitude $M_{xy}$. a) Top surface, b) Back surface.
@@
The contrast of the peaks in an ODMR graph can provide important information on the quality of the color centers. It not only regards the signal-to-noise ratio (SNR), but is also affected by effects of decoherence and inhomogeneous broadening, which significantly impact the optical properties of the emitters. These phenomena can be caused by many factors, including the presence of strain fields which are investigated here. Figure 6 demonstrates how contrast is exacerbated close to the diamond-SiO2/Si interface. The initial ODMR intensity contrast at the top surface is approximately 5.32%, going up to an average of 5.51% in the region where the NV centers exhibit the best signal, and finally dropping down to approximately 5.15% close to the interface.
The average Full Width Half Maximum (FWHM) of the peaks in the ODMR spectra follows a similar trend, as it seems to diminish close to the interface. It drops from an average of $6.08$ MHz close to the top surface, down to an average of $5.70$ MHz close to the interface. This is an unexpected result, as one would anticipate that the observed linewidth worsens under increased strain conditions, similar to how contrast is exacerbated.
Regarding the experimental setup, problems occasionally arise when focusing on the back surface of a thick substrate, due to limits set by the working distance of the objective lens, as well as obstruction by the microwave wire. This creates the need for thinning of the diamond layer by O2 plasma etching; a rather time-consuming process. Last but not least, the ODMR measurement process itself is another factor to take into consideration concerning the quality of the measurements. Long scanning durations make the samples more prone to thermal drifts and can introduce experimental parameter fluctuations. The time inefficiency is further exacerbated by the long direct bonding processing time of approximately a week. It is worth noting that future DOI substrates may be comprised of diamonds with a thickness of a few micrometers or even in the nanometer range. In such cases, the aforementioned strain effects and the quality of the emitters are bound to differ and be more localized.

### v2

@@ 1 Introduction
One of the challenges in achieving robust quantum communication and high-fidelity quantum computing has long been the entanglement generation rate between NV centers, which is limited by the relatively low fraction of emitted coherent photons that can be detected. A solution to address this problem is to integrate NV centers into optical cavities with a high ratio of quality factor to mode volume. This is done to increase the zero-phonon line (ZPL) emission and improve collection efficiency [1].
@@
However, despite the effectiveness of direct bonding and subsequent annealing in enhancing interface quality, this fabrication process can introduce unintended strain effects within the diamond structure. It is probable that the diamond, having undergone the aforementioned chemical processing, will exhibit further stress effects on the surface which is bonded to the insulator substrate. Furthermore, part of the direct bonding process takes place under a temperature of 200°C. During the cooling period, each of the bonded materials contracts at different rates, according to their unique thermal expansion coefficients. This mismatch between the thermal expansion rates between the diamond and the silicon dioxide/silicon can lead to the generation of strain at their interface [14, 15], and potentially further deep into the diamond lattice.
These strain effects can cause deformations to the crystal lattice, which result in inhomogeneous broadening and degradation of the spin dephasing time of the NV centers. Therefore, it is deemed critical to be able to understand and visually map the effects of strain, providing meaningful insight into the quality of the bonded samples.
@@ 2 Direct Bonding Process Overview
The fabrication of the diamond-on-insulator (DOI) substrate utilized in this work follows a hydrophilic direct bonding process, outlined in Figure 1. No adhesives were used in the bonding, which relied entirely on surface chemical activation and low-temperature annealing to achieve interfacial adhesion.
Figure 1: Schematic overview of the hydrophilic direct bonding process used to fabricate the diamond-on-insulator (DOI) substrate. The process begins with a type Ib, single-crystal, double-side polished, (100)-oriented diamond substrate, and a Si substrate, coated with a 300 nm PECVD SiO2 layer, and diced into 25$\times$25 mm2 chips. Both substrates undergo chemical activation. The diamond is cleaned using Piranha solution (3:1 H2SO4:H2O2) at 75C for 30 min and rinsed in deionized (DI) water for 5–10 min. Meanwhile, the SiO2/Si substrate is rinsed with DI water for 5 min, and is then subjected to O2 plasma treatment (1000 W, 5 min, with 400 sccm O2 flow) using a PVA Tepla system. These steps remove surface contaminants and enrich the surface with silanol (Si–OH) groups, promoting hydrophilicity. After cleaning, the substrates are brought into contact at room temperature (20C) and ambient humidity ($\sim$40%), allowing van der Waals and hydrogen bonding to form across a thin interfacial water layer. This pre-bonding interaction is stabilized through a low-temperature anneal at 200C for 24 h, under N2 flow. The annealing induces interfacial dehydration, resulting in covalent bond formation, thus completing the DOI heterostructure. The bonded stack typically exhibits high mechanical integrity with no visible voids or delamination.
@@
Hydrophilic direct bonding is a process that involves creating a bond between surfaces treated to exhibit hydrophilic properties. The first step is surface activation, which prepares the surfaces for subsequent bonding by introducing reactive hydroxyl groups (–OH). For the diamond substrate, this involves thorough cleaning using a Piranha solution, a highly oxidative mixture of sulfuric acid (H2SO4) and hydrogen peroxide (H2O2) in a 3:1 ratio. The Piranha solution does not etch the diamond, but removes organic contamination and promotes surface hydrophilicity. This process results in the termination of the diamond surface with hydroxyl groups (C–OH), enhancing its reactivity, by providing essential reactive sites for bonding with other molecules and facilitating further chemical reactions [13, 16].
The insulator substrate is composed of a 300 nm, PECVD-grown layer of SiO2 on Si. Hydrophilic bonding relies on the formation of silanol (Si–OH) groups on the SiO2 surface. Therefore, following rinsing of the SiO2/Si substrate with DI water, it undergoes O2 plasma treatment, a common method for cleaning and oxidizing silicon surfaces [13]. This treatment both removes surface contamination and increases silanol group density by oxidizing siloxane (Si–O–Si) bridges in the presence of water vapor. These hydroxyl-terminated surfaces are hydrophilic and attract adsorbed water molecules from the environment.
Figure 2: Illustration of the interfacial bonding mechanism between diamond and SiO2. Increased surface roughness enhances the density of hydroxyl (–OH) terminations, facilitating more effective hydrogen bonding and improving the overall bonding success. Reproduced with permission from Chen et al.[11]. © 2025 AIP Publishing.
Once activated, both the diamond and SiO2/Si surfaces become highly hydrophilic due to the formation of surface hydroxyl (–OH) groups. This promotes the adsorption of ambient water molecules, leading to the formation of a thin hydration layer on each surface. When the substrates are brought into contact under ambient conditions, hydrogen bonding and van der Waals interactions across this interfacial water layer enable initial adhesion and stabilize the alignment of the surfaces (see Figure 2). Specifically, hydroxyl groups such as C–OH and Si–OH participate in hydrogen bonding, which, although individually weak, collectively contribute significantly to the mechanical integrity of the interface[17].
@@
Following initial contact, the DOI stack undergoes annealing, with temperature and duration serving as critical parameters for achieving robust interfacial bonding. Utilizing low-temperature annealing at 200C, the bond strength is enhanced through dehydration reactions between adjacent silanol groups, producing strong Si–O–C covalent bridges across the interface[17]. This irreversible covalent bonding process completes the formation of the DOI heterostructure, as expressed by the reaction seen below:
@@
C-OH + HO-Si -¿ C-O-Si + H2O
Throughout this process, care was taken to maintain surface roughness within an optimal range. The surface roughness values of the diamond and the SiO2/Si substrate fall within the 2–5 nm range typically reported as suitable for reliable hydrophilic bonding[11]. Notably, no high-temperature processing or aggressive etching was applied that could significantly modify surface topography prior to bonding.
The bonded interface remains intact through post-processing steps, and no delamination or void formation is typically observed. This is consistent with the formation of strong interfacial bonds and the maintenance of a mechanically coherent interface. The process flow has been replicated multiple times, further validating its general applicability and reproducibility.
3 Strain Effects in the Diamond-on-Insulator System
3.1 Origin and Mechanisms of Strain in Directly Bonded DOI Structures
The direct bonding process of a diamond-on-insulator (DOI) substrate inevitably induces strain at the interface, with the most significant contribution arising from thermal expansion mismatch between the bonded materials[15], since prior to bonding, the residual stress in both the diamond and SiO2/Si substrates is minimal[18]. The diamond used did not undergo post-growth annealing or epitaxy, which would otherwise introduce non-uniform stress[19]. Additionally, optical inspection revealed no bowing, indicating negligible macroscopic stress. Similarly, the underlying substrate remains mechanically relaxed[18]. Any pre-existing uniform residual stress is accounted for as a constant offset in the ODMR measurements, as the observed strain is induced post-bonding.
The annealing process at 200C, necessary for bond stabilization, introduces thermal strain in the DOI heterostructure upon cooling to room temperature. This strain arises due to the mismatch in thermal expansion coefficients between the constituent materials. Diamond exhibits a low thermal expansion coefficient, with $\alpha_{\mathrm{D}}\approx 1.1\times 10^{-6}$ K-1 at room temperature and $\alpha_{\mathrm{D}}\approx 2.3\times 10^{-6}$ K-1 at 200C[20]. In contrast, amorphous silica (SiO2) has an expansion coefficient of $\alpha_{\mathrm{SiO_{2}}}\approx 2.5\times 10^{-6}$ K-1 at room temperature[21] and $\alpha_{\mathrm{SiO_{2}}}\approx 3.0\times 10^{-6}$ K-1 at 200C[21, 22]. Silicon, on the other hand, expands significantly more than the diamond, with $\alpha_{\mathrm{Si}}\sim 2.6\times 10^{-6}$ K-1 at room temperature and $\alpha_{\mathrm{Si}}\sim 3.3\times 10^{-6}$ K-1 at 200C[23, 24]. Table 1 summarizes the relevant thermal expansion coefficients. Since the SiO2 layer is only 300 nm thick, while the underlying Si substrate is several hundred microns thick, the thermal and mechanical behavior of the bonded insulator stack is dominated by the silicon. The diamond layer, in turn, is constrained by this underlying Si substrate across the interfacial SiO2 layer, which behaves elastically under these conditions. Upon cooling, the Si substrate contracts more than the diamond, imposing compressive in-plane strain on the diamond layer[25]. For a temperature drop of $\Delta T\sim 180$ K, the resulting in-plane thermal mismatch strain is estimated as $\varepsilon\sim(\alpha_{\mathrm{Si}}^{200^{\circ}\mathrm{C}}-\alpha_{\mathrm{D}}^{200^{\circ}\mathrm{C}})\Delta T\approx-1.8\times 10^{-4}$.
Table 1: Linear thermal expansion coefficients ($\alpha$) of materials in the DOI stack.
Material $\alpha$ (RT) (×$\times 10^{-6}$ K-1) $\alpha$ (200C) (×$\times 10^{-6}$ K-1)
Diamond 1.1$\sim$ 1.1∼ 1.1 2.3$\sim$ 2.3∼ 2.3
PECVD SiO2 2.5$\sim$ 2.5∼ 2.5 3.0$\sim$ 3.0∼ 3.0
Silicon 2.6$\sim$ 2.6∼ 2.6 3.3$\sim$ 3.3∼ 3.3
While the SiO2 layer serves as the direct interface between diamond and Si, it is important to consider whether it could act as a strain-relieving buffer. In particular, viscous relaxation or interfacial slippage could potentially mitigate strain transfer. However, previous studies report that viscous flow in SiO2 requires temperatures above 900–950C[26], well beyond the 200C annealing conditions used here. At these lower temperatures, the oxide behaves elastically. Moreover, the formation of covalent interfacial bonds effectively suppresses slippage at the diamond–SiO2 interface[11]. As a result, the bonded stack behaves as a mechanically coherent system, with strain transmitted directly from the silicon substrate to the diamond through the thin, elastically deforming oxide layer.
Surface roughness is another potential contributor to strain. However, in our study, the bonded diamond surface had an initial roughness of $R_{a}<2$ nm, preserved through the Piranha and DI water cleaning steps. O2 plasma etching was applied to the unbonded top diamond surface, with no measurable change in surface roughness. The SiO2 layer had a measured RMS roughness of $S_{q}=4.44$ nm prior to bonding. Both these values fall within the 2–5 nm range considered suitable for uniform bonding[11].
Finally, localized surface defects, such as scratches or implantation damage, are known to produce sharp strain gradients and localized ODMR shifts[27]. In contrast, the strain fields observed in our bonded samples are spatially extended and homogeneous over tens of microns, without any localized stress anomalies. Delamination is also not typically observed, following this direct bonding process. These features are inconsistent with surface-defect-induced strain, and instead match the expected thermomechanical strain distribution in a bonded multilayer stack. Therefore, the strain analyzed in this work originates predominantly from thermal expansion mismatch, transmitted through the SiO2 layer without significant relaxation.
@@ 3.2.1 Spin-Strain Interaction
In reality, numerous imperfections exist within the diamond lattice, making the spin-strain interaction non-negligible. According to Udvarhelyi et al. (2018) [29], the spin-strain interaction for the ground-state spin triplet can be described as a deformation in the diamond crystal, causing the strain tensor $\varepsilon_{ij}$ to couple to the NV spin through the symmetry-allowed Hamiltonian:
@@
[…]
@@
Table 2: Spin-strain ($h$) coupling-strength parameters calculated from density functional theory by Udvarhelyi et al. (2018) [29].
@@ 3.2.2 Eigenfrequencies Including the Zeeman Term
$\frac{\hat{H}_{B}}{h}=\gamma\mathit{B}\cdot\mathbf{S},$ (5)
where $\gamma$ is the electron gyromagnetic ratio, 2.82.82.8 MHz/G. After diagonalizing this matrix and including the spin-strain interaction Hamiltonian, the following equation is obtained:
@@
where $M_{x}$ and $M_{y}$ represent the shear (transverse) strain components, corresponding to lattice distortions perpendicular to the NV axis. These transverse components primarily influence the magnitude of peak splitting in the ODMR spectra. Therefore, the effects of strain can be directly characterized by ODMR [30], as lattice distortions lead to measurable frequency shifts and splittings of the spin resonance transitions.
@@ 4 Experimental Section
The primary objective of this study is to measure the effects of strain in diamonds that have undergone direct bonding with an insulator ($\mathrm{SiO_{2}/Si}$) substrate. To achieve this, NV centers are used due to their excellent properties that make them highly sensitive to external magnetic fields and strain fields. As already described above, strain can cause splitting and shifts in the energy levels of NV centers. To enable these measurements, a confocal Optically Detected Magnetic Resonance (ODMR) setup is used, as seen in Figure 3. With ODMR, these changes in the energy states and spins can be quantified in the form of the presence of magnetic, electric, strain, or temperature variations. The frequency shift $\Delta f_{\mathrm{shift}}$ and frequency splitting $\Delta f_{\mathrm{split}}$ are related to the strain tensor components $\epsilon_{\mathrm{ij}}$, through Equation 3.
Figure 3: Schematic illustration of the confocal Optically Detected Magnetic Resonance (ODMR) setup used in this work. A 532 nm CW-laser provides optical excitation of the NV centers, with fluorescence collected via a high-NA objective and directed through a dichroic mirror and filters to a single-photon counting module (SPCM). Microwave excitation is delivered via a wire loop placed near the diamond surface. A bias magnetic field ($B_{z}$) is applied along the NV axis to lift spin degeneracy, enabling detection of frequency shifts and splittings associated with strain.
@@
For the experiment, a $10\times 10\,\mathrm{mm^{2}}$ bulk diamond substrate (supplied by EDP), with a thickness of 300$\sim$ 300∼ 300 µm is investigated, hereafter referred to as $\mathrm{DS_{PB}}$. It is a CVD-grown substrate, cut along the {100} crystallographic plane, that contains randomly distributed ensembles of NV centers with a concentration of roughly 15 ppb. The diamond substrate was double-side polished, exhibiting an initial surface roughness of $R_{a}<2$ nm, as specified by the supplier. The SiO2 surface on the Si wafer, prior to bonding, exhibited a root mean square (RMS) roughness of $S_{q}=4.44$ nm, as measured by AFM.
[…]
It should be noted that the following measurements are performed in the presence of an external bias magnetic field $B_{z}=15.7$ G (or 43.9 MHz splitting) along the $z$-axis. A 532 nm green laser is used for optical excitation, at a power of 0.25 mW, and the emitted fluorescence is collected via a single-photon counting module (COUNT-10C-FC), coupled to a fiber.
@@ 5 Results and Discussion
Figure 6: ODMR 2D scans performed in a $\times 60$ 60 µm2 area, depicting the effects of the volumetric strain amplitude $M_{z}$. a) Top surface, b) Back surface.
[…]
The milder volumetric strain effects at the top surface are attributed to intrinsic strain in the diamond. However, moving closer towards the back surface shows increased strain effects, which are assumed to have been caused by the thermal expansion mismatch that occurs during the direct bonding process.
Figure 7: ODMR 2D scans performed in a $\times 60$ 60 µm2 area, depicting the effects of the transverse strain amplitude $M_{xy}$. a) Top surface, b) Back surface.
@@
The contrast of the peaks in an ODMR graph can provide important information on the quality of the color centers. It not only regards the signal-to-noise ratio (SNR), but is also affected by effects of decoherence and inhomogeneous broadening, which significantly impact the optical properties of the emitters. These phenomena can be caused by many factors, including the presence of strain fields which are investigated here. Figure 9 demonstrates how contrast deteriorates close to the diamond-SiO2/Si interface. The initial ODMR intensity contrast at the top surface is approximately 5.32%, increasing to an average of 5.51% in the region where the NV centers exhibit the best signal, and finally decreasing to approximately 5.15% close to the interface.
The average Full Width Half Maximum (FWHM) of the peaks in the ODMR spectra follows a similar trend, as it seems to diminish close to the interface. It drops from an average of 6.086.086.08 MHz close to the top surface, down to an average of 5.705.705.70 MHz close to the interface. This is an unexpected result, as one would anticipate that the observed linewidth worsens under increased strain conditions, similar to how contrast decreases.
Regarding the experimental setup, problems occasionally arise when focusing on the back surface of a thick substrate, due to limits set by the working distance of the objective lens, as well as obstruction by the microwave delivery wire. This necessitates thinning of the diamond layer by O2 plasma etching; a rather time-consuming process. Lastly, the ODMR measurement process itself is another factor to take into consideration concerning the quality of the measurements. Long scanning durations make the samples more prone to thermal drifts and can introduce experimental parameter fluctuations. The time inefficiency is further exacerbated by the long direct bonding processing time of approximately a week. It is worth noting that future DOI substrates may be comprised of diamonds with a thickness of a few micrometers or even in the nanometer range. In such cases, the aforementioned strain effects and the quality of the emitters are bound to differ and be more localized.